Grade 12Physical Chemistry

what is the realtion between n factor and normality.

Ankit Saini
8 Years agoGrade 12

2 Answers

Deepak Kumar Shringi
8 Years ago
ormality (N)= Number of gm. equivalents of solute/volume of solution( liter)

Number of gm. equivalents of solute = w/E =weight of solute/Equivalent weight of solute

N = w / EV(l)

N/10 ( decinormal )= When 1/10 gm.equivalents of solute are present in one liter of solution,then solution is decinormal.

N/100 ( centi normal )= When 1/100 gm.equivalents of solute are present in one liter of solution,then solution is centinormal.

N/1000 ( milli normal )= When 1/1000 gm.equivalents of solute are present in one liter of solution,then solution is milli norma
Subhash Kumar Mehta
8 Years ago
Required relation is N=F*M where N=normality, F=factor and M=molarity »Thanks »Regards »Subhash Kumar »St. Xavier`s College, Ranchi